“Blast it! said David Wilson, president of Teledex Company. “Weve just lost the bid on the Koopers job by $3,000. It seems5. At the end of the year, the company assembled the following actual cost data relating to all jobs worked on during the yeab. Compute the underapplied or overapplied overhead for the year, assuming that departmental overhead rates are used. (Enter
Answer #1
Fabricating Machining Assembly Total Plant
Direct Labor 201000 100500 301500 603000
Manufacturing Overhead 351750 402000 90450 844200
Overhead Rate 175% 400% 30% 140%
a.Platwide predetermined rate 140% of Direct labor cost
Overhead applied                   816,200
Actual Overhead                   867,000
Underapplied                     50,800
Overhead applied Actual Overhead Difference
Fabricating                   369,250                     361,000         8,250 Overapplied
Machining                   436,000                     421,000       15,000 Overapplied
Assembly                     78,900                       85,000       (6,100) Under applied
Total Plant                   884,150                     867,000       17,150 Overapplied

B Fabricating 201000 351750 -R =B3/B2 /D Machining 100500 402000 =C3/C2 DE Assembly Total Plant 301500 =B2+C2+D2 90450 =B3+C3